OpenStudy (anonymous):

you drove to your cousins house which is 460 miles away. after 2 hours you had goe 100 miles. after 8 hours you reached your destination. write an equation that gives the number of miles you had driven y in terms of the number of hours you had driven.

OpenStudy (anonymous):

Hello there

OpenStudy (anonymous):

How are you?

OpenStudy (anonymous):

Please help me...

OpenStudy (anonymous):

Hello, I'm doing fine and trying to think of a way to solve this easily. Don't suppose this is multiple choice is it?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

this a hmwrk problem...

OpenStudy (anonymous):

i am in 7th grade...

OpenStudy (anonymous):

Alright then give me a sec to see how to write this out.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Still here, are you doing Linear functions or...what type of math?

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Linear Functions is apart of algebra, you said this is homework so what type of stuff were you learning? Like formulas that could help you plug this in.

OpenStudy (anonymous):

slope...we have to write it in slope intercept form.... no formulas my teacher told me...

OpenStudy (anonymous):

Okay then well our formula for slope is m = (y2 - y1) / (x2 - x1), familiar yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good. Now our points we'll be working with is (2,100) and (8,460). Now we need to plug them into the formula, can you do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how did you get the pionts?

OpenStudy (anonymous):

Because we drove 100 miles after 2 hours and in 8 hours we reached our total 460 miles. So think of it as x=time and y=distance.

OpenStudy (anonymous):

oh... thank you so much... you want me to do it right now?

OpenStudy (anonymous):

Yep, now plug in our points and show me how the equation looks afterward. :)

OpenStudy (anonymous):

ok wait...

OpenStudy (anonymous):

i got 6

OpenStudy (anonymous):

That's just part of our solution: m=(460-100)/(8 - 2) =360/ 6. Do you understand that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do we do next?

OpenStudy (anonymous):

Alright now we know that gives us 60. What we need is the linear function, our equation for this is y=m*t + b. Right now we have b = y - 60*t. m is slope, t is time, y is distance and b is our y-intercept.

OpenStudy (anonymous):

but i got 6 how did you get 60?

OpenStudy (anonymous):

Because our slope equation was m=(460-100)/(8 - 2) =360/ 6. 360 divided by 6=60

OpenStudy (anonymous):

oh yeah...sorry

OpenStudy (anonymous):

It's okay, you understand the rest tho right?

OpenStudy (anonymous):

yes...thank you so much!

OpenStudy (anonymous):

Lol no prob so you need me to walk you through the rest or...?

OpenStudy (anonymous):

no i got it...

OpenStudy (anonymous):

You r the BEST!!

OpenStudy (anonymous):

Lol just doing my job. Take care. :)

OpenStudy (anonymous):

you too!!

OpenStudy (anonymous):

do you have a email...we can keep in touch...

OpenStudy (anonymous):

My email gets flooded x.x Lol but I'm on here alot so you could reach me through this sites mailing system easier . :)

OpenStudy (anonymous):

what do you mean by flooded?

OpenStudy (anonymous):

Tons of ppl emailing allll day. And I don't check it as much

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

so i will see ya later.!

OpenStudy (anonymous):

Yep bye.